OpenStudy (calculator):

EXPAND sqrt{/frac{1-x}{1+x}}

OpenStudy (calculator):

\[\Huge \sqrt{\frac{1-x}{1+x}}\]

OpenStudy (calculator):

up to power x^2

OpenStudy (calculator):

i get \[1-x+\frac{x^2}{2} - \frac{x^3}{8} - \frac{3x^4}{64}\]

OpenStudy (calculator):

@saifoo.khan

OpenStudy (calculator):

wolfram got \[-\frac{3 x^5}{8}+\frac{3 x^4}{8}-\frac{x^3}{2}+\frac{x^2}{2}-x+1\]

OpenStudy (saifoo.khan):

Unfortunately i forgot the formula! D:

OpenStudy (anonymous):

Is it a maclaurin series?

OpenStudy (calculator):

you can try that

OpenStudy (anonymous):

I want to know what is meant by Expand, since that's the only thing that comes to mind.

OpenStudy (calculator):

you can use this \[(x+a)^n=\sum ^{n}_{ k=0} \left(\begin{matrix}n \\ k\end{matrix}\right) x^k a^{(n-k)}\]

OpenStudy (calculator):

or \[(1+x)^n = 1+nx+\frac{(n)(n-1)}{2!}(x^2)+...\]

OpenStudy (anonymous):

That works for \(n\in \mathbb{N}\).

OpenStudy (anonymous):

In this case you have a square root...

OpenStudy (calculator):

still can, watch

OpenStudy (calculator):

\[\sqrt{\frac{1-x}{1+x}} \\ \\ \frac{\sqrt{1-x}}{\sqrt{1+x}} \\ \\ (1-x)^\frac{1}{2}(1+x)^{-\frac{1}{2}}\]

OpenStudy (calculator):

expand separately

OpenStudy (calculator):

mix them by multiplying together

OpenStudy (anonymous):

Yeah, my problem is with \(1/2 \notin \mathbb{N}\)

OpenStudy (calculator):

I don't think it should be natural number

OpenStudy (calculator):

i mean real number will do

OpenStudy (anonymous):

How can you have summations or binomial coefficients with rational numbers?

OpenStudy (amoodarya):

interval of x is ?

OpenStudy (calculator):

the formula list says can use

OpenStudy (calculator):

we use the second one (1+x)^n

OpenStudy (calculator):

http://en.wikipedia.org/wiki/Binomial_series look at the History section

OpenStudy (amoodarya):



OpenStudy (amoodarya):



OpenStudy (calculator):

or maybe its \[-\frac{3 x^5}{8}+\frac{3 x^4}{8}-\frac{x^3}{2}+\frac{x^2}{2}-x+1\]

OpenStudy (calculator):

hmm

OpenStudy (calculator):

based from the formula as long as its in the from ("1"+x)^n should be fine to expand, i tried it

OpenStudy (calculator):

from \[(1-x)^{1/2} \\ \\ n=-1/2, x=-x(for formula) \\ \\ [1-\frac{x}{2}+\frac{(\frac{1}{2})(\frac{1}{2}-1)}{2}(x^2)] \\ \\ 1-\frac{x}{2}-\frac{x^2}{8}\]